OpenStudy (sw050399):

The base of a solid is a circle of radius a, and its vertical cross sections are equilateral triangles. Find the radius of the circle if the volume of the solid is 10 cubic meters.

mathslover (mathslover):

I think calculus would be needed here

OpenStudy (sw050399):

You are correct. This is a practice problem from my Calculus I class that I didn't really understand. If you have anything to contribute, it would be greatly appreciated :)

mathslover (mathslover):

it is clear that circle is in the XY plane. So the integration would be along X or Y axis. Well it does not matter.

OpenStudy (sw050399):

Okay, I get that part.

mathslover (mathslover):

Well you can put the triangles perpendicular to the X-axis

mathslover (mathslover):

I am going to use x - axis here.

OpenStudy (sw050399):

Not a problem.

mathslover (mathslover):

Let dA be the area of the triangles . So volume = \(\int {dA dx}\)

mathslover (mathslover):

calculate dA , \(dA = \cfrac{1}{2} b\times h\)

OpenStudy (sw050399):

Because it's an equilateral triangle, would you have to involve \[((s^2)(\sqrt{3}))/4 \] somehow?

mathslover (mathslover):

well .. do you agree that base = 2y

OpenStudy (sw050399):

Yes. It's an equilateral triangle and because it's perpendicular to the x axis all sides are equal to 2y.

mathslover (mathslover):

Yep! \(h = y \tan 60\) or \(h = y\sqrt{2}\)

mathslover (mathslover):

So \(dA = ?\) can u tell me?

mathslover (mathslover):

You have to just apply : 1/2 bh formula.. you have b = 2y and h = y\(\sqrt{2}\)

OpenStudy (sw050399):

How did you get sqrt2? Isn't it sqrt3?

mathslover (mathslover):

h = y tan 60 ... agreed?

mathslover (mathslover):

Yeah!

mathslover (mathslover):

sqrt{3} sorry

mathslover (mathslover):

So we have \(dA = ?\) ..

OpenStudy (sw050399):

\[dA = (1/2)(2y)(y \sqrt{3})\] \[= y^2 \sqrt{3}\]

mathslover (mathslover):

yes!

mathslover (mathslover):

Find \(\int y^2 \sqrt{3} dx\)

mathslover (mathslover):

you can shorten your work : \(y^2 = r^2 -x^2\) put that there on the place of \(y^2\)

OpenStudy (sw050399):

\[\int\limits(r^2 - x^2)(\sqrt{3})dx\] \[((r^3)/3 - (x^3)/3)*((6\sqrt{3})/3)\] \[(1/3)(r^3 - x^3)(2\sqrt{3}) + C\]

OpenStudy (sw050399):

I forgot the + C in the 2nd line...

mathslover (mathslover):

I think it will be : \(\sqrt{3}(r^2 x - \cfrac{x^3}{3}) + C\)

mathslover (mathslover):

check ur method again!

mathslover (mathslover):

\(\int (r^2-x^2)\sqrt{3} dx\) \(\sqrt{3} \int (r^2-x^2)dx\) \(\sqrt{3} [\int r^2 dx - \int x^2 dx]\) \(\sqrt{3} [r^2x - \cfrac{x^3}{3}]+C\) got it?

OpenStudy (sw050399):

Yes. I actually got right before you posted the work... I just forgot that we were integrating with respect to x.

mathslover (mathslover):

Now evaluate it over the range : -r to r ...

mathslover (mathslover):

note: evalutate it in X

mathslover (mathslover):

then equate it to 10 unit^3 find r

OpenStudy (sw050399):

\[r = \sqrt{30/4\sqrt{3}} - C\]

mathslover (mathslover):

would it be cube root?

OpenStudy (sw050399):

Yes, I just forgot to type that part.

mathslover (mathslover):

ok! and I think there would be no -C

OpenStudy (sw050399):

Right, because the the other part is just a constant and adding some some that stays the same all the time isn't going to do anything.

mathslover (mathslover):

So you have your answer .. solve that by using calculator or let it be like that only.

mathslover (mathslover):

though @hartnn would check out the whole method ..

OpenStudy (sw050399):

Thank you for your help!! I'll try that answer and see how it goes. I think we're right, but my teacher's kinda crazy, so I'll let you know. Thank you again!! :)

mathslover (mathslover):

You're welcome and Best of Luck!